# Cost-Sensitive Learning to Defer to Multiple Experts with Workload Constraints

### v2

@@ Abstract
Learning to defer (L2D) aims to improve human-AI collaboration systems by learning how to defer decisions to humans when they are more likely to be correct than an ML classifier. Existing research in L2D overlooks key aspects of real-world systems that impede its practical adoption, namely: i) neglecting cost-sensitive scenarios, where type 1 and type 2 errors have different costs; ii) requiring concurrent human predictions for every instance of the training dataset and iii) not dealing with human work capacity constraints. To address these issues, we propose the deferral under cost and capacity constraints framework (DeCCaF). DeCCaF is a novel L2D approach, employing supervised learning to model the probability of human error under less restrictive data requirements (only one expert prediction per instance) and using constraint programming to globally minimize the error cost subject to workload limitations. We test DeCCaF in a series of cost-sensitive fraud detection scenarios with different teams of 9 synthetic fraud analysts, with individual work capacity constraints. The results demonstrate that our approach performs significantly better than the baselines in a wide array of scenarios, achieving an average $8.4\%$ reduction in the misclassification cost. The code used for the experiments is available at https://anonymous.4open.science/r/deccaf-1245/
@@
1 Introduction
An increasing body of recent research has been dedicated to human-AI collaboration (HAIC), with several authors arguing that humans have complementary sets of strengths and weaknesses to those of AI [De-Arteaga et al., 2020, Dellermann et al., 2019]. Collaborative systems have demonstrated that humans are able to rectify model predictions in specific instances [De-Arteaga et al., 2020], and have shown that humans collaborating with an ML model may achieve synergistic performance: a higher performance than humans or models alone [Inkpen et al., 2022]. In high-stakes scenarios where ML models can outperform humans, such as healthcare [Gulshan et al., 2016], HAIC systems can help address safety concerns (e.g., the effect of changes in the data distribution [Gama et al., 2014]), by ensuring the involvement of humans in the decision-making process.
The state-of-the-art framework to manage assignments in HAIC is learning to defer (L2D) [Charusaie et al., 2022, Hemmer et al., 2022, Raghu et al., 2019b, a, Mozannar and Sontag, 2020b, Mozannar et al., 2023, Madras et al., 2018b, Steege, 2023, Verma and Nalisnick, 2022a, Verma et al., 2023]. L2D aims to improve upon previous approaches, such as rejection learning [Chow, 1970, Cortes et al., 2016], which defer based solely on the ML model’s confidence, by also estimating the human confidence in a given prediction and passing the instance to the decision-maker who is most likely to make the correct decision.
Previous work in L2D does not address several key aspects of collaborative systems [Leitão et al., 2022]. In real-world scenarios, multiple human experts are employed to carry out the classification task, as the volume of instances to process cannot be handled by a single human. However, only a small subset of L2D research focuses on the multi-expert setting, where decisions have to be distributed throughout a team comprised of a single ML classifier and multiple human experts [Keswani et al., 2021, Hemmer et al., 2022, Verma et al., 2023]. To the best of our knowledge, Verma et al. [2023] propose the only two consistent multi-expert L2D formulations, with both approaches assuming the existence of every expert’s predictions for all training instances. In real-world applications, we often have more limited data availability, with only a single expert [De-Arteaga et al., 2020] or a small subset of the experts [Gulshan et al., 2016] providing predictions for each instance. This means that a practitioner aiming to train L2D algorithms would have to purposefully collect the set of every expert’s predictions for a sample of instances, which could incur large costs or even be unfeasible [Leitão et al., 2022]. We will propose L2D architectures that allow for assigners to be trained assuming that each instance is accompanied by the prediction of only one expert out of the team.
Current multi-expert L2D methods also neglect human capacity limitations [Leitão et al., 2022]. Should there be an expert that consistently outperforms the model and all the other experts, the optimal assignment would be to defer all cases to that expert, which in practice is not feasible. Furthermore, current work neglects cost-sensitive scenarios, where the cost of misclassification can be class or even instance-dependent (e.g., in medicine, false alarms are generally considered less harmful than failing to diagnose a disease).
To address the aforementioned L2D limitations, we propose the deferral under cost and capacity constraints framework (DeCCaF): a novel deferral approach to manage assignments in cost-sensitive human-AI decision-making, which respects human capacity constraints. Our method is comprised of three separate parts, represented schematically in Figure 1: (1) an ML classifier modelling the probability of the target class given the instance features; (2) a human expertise model (HEM) that models the probabilities of correctness of each of the experts in the team; and (3) an assigner that computes the best possible set of assignments given misclassification costs and capacity constraints.
Due to the lack of sizeable datasets with multiple human predictions, and the high costs associated with producing one, we empirically validate DeCCaF in a series of realistic cost-sensitive fraud detection scenarios, where a team of 9 synthetic fraud analysts and one ML classifier are tasked with reviewing financial fraud alerts. We conclude that, across all scenarios, DeCCaF performs similarly to, or significantly better than L2D baselines, as measured by average misclassification costs.
@@ 1 Introduction
In Section 2, we describe the relevant related work, focusing on recent developments in multi-expert L2D and examining the shortcomings of the deferral systems proposed so far. We also discuss current practices in L2D evaluation, particularly the use of synthetically generated expert predictions, and how they can be improved. We then describe DeCCaF in Section 3, first by formulating a novel training method, compatible with limited data availability, for the commonly used classifier-rejector framework [Mozannar and Sontag, 2020a, Verma and Nalisnick, 2022a, Verma et al., 2023, Cortes et al., 2016], demonstrating that it produces optimal unconstrained deferral. As the traditional classifier-rejector approach fails to consider the existence of human work capacity limitations, we propose a novel formulation for global assignment optimization under capacity constraints. In Section 4 we detail a realistic fraud detection experimental setup as well as the method for the synthetic expert prediction generation. We also discuss the capacity-aware baselines used in our L2D benchmark. The experimental results are displayed and analyzed in Section 5.
@@ 2.1 Current L2D Methods
The simplest deferral approach in the literature is rejection learning (ReL), which dates back to the work of Chow [1970], Cortes et al. [2016]. In a HAIC setting, ReL defers to humans the instances that the model rejects to predict. [Madras et al., 2018b, Raghu et al., 2019a]. A simple example [Hendrycks and Gimpel, 2017] is to obtain uncertainty estimates of the model prediction for each instance, rejecting to predict if the uncertainty estimate is above a given threshold.
@@
Madras et al. [2018b] criticize ReL, arguing that it does not consider the performance of the human involved in the task and propose learning to defer (L2D), where the classifier and assignment system are jointly trained, taking into account a single model and a single human and accounting for human error in the training loss. Many authors have since contributed to the single-expert framework: Mozannar and Sontag [2020b] show that the loss proposed by Madras et al. [2018b] is inconsistent, proposing a consistent surrogate loss that yields better results in testing; Verma and Nalisnick [2022b] critique the approach of Mozannar and Sontag [2020b], demonstrating that their surrogate loss has a degenerate parameterization, causing miscalibration of the estimated probability of expert correctness. Keswani et al. [2021] observe that decisions can often be deferred to one or more humans out of a team, expanding L2D to the multi-expert setting. Verma et al. [2023] propose the first consistent and calibrated surrogate losses for the multi-expert setting, by adapting the work of Verma and Nalisnick [2022b] and the softmax surrogate loss of Mozannar and Sontag [2020b]. All aforementioned studies focus on deriving surrogates for the $0-1$ loss, meaning they are not directly applicable to cost-sensitive scenarios, where the cost of erring can be class-dependent (i.e., different costs for false positive and false negative errors), or even instance-specific (i.e., every instance has an associated misclassification cost).
@@
Another key facet of L2D research is the interplay between the two components of an assignment system: the rejector (which decides whether to defer, and to whom) and the classifier (which produces automatic predictions if the rejector chooses not to defer). Mozannar and Sontag [2020b] argue that the main classifier should specialize on the instances that will be assigned to it, in detriment of those that will not. This is done by jointly training the classifier and the rejector, without penalizing the classifier’s mistakes on instances that the rejector defers to an expert. The approach of Verma et al. [2023] differs in that the rejector and the classifier are trained independently, meaning that the classifier is encouraged to predict correctly on all instances, likely to be deferred or not. While Mozannar et al. [2023] demonstrate that joint learning enables the system to theoretically achieve an improvement in overall performance, the one-vs.-all approach of Verma et al. [2023] outperforms their adaptation of the softmax loss of Mozannar and Sontag [2020a], due to the same calibration problems observed in the single-expert setting. Furthermore, joint learning may not be suitable for real-world applications, as, by design, the instances that are most likely to be deferred are those in which the classifier will perform worse, which would make the system highly susceptible to changes in human availability, should the AI have to predict on instances that were originally meant to be deferred.
@@
As previously mentioned, the key barrier to the adoption of L2D methods is that they require predictions from every human in the team, for every training instance. Current L2D research tackling data-efficient learning focuses only on the single-expert setting, using active-learning principles [Charusaie et al., 2022] to construct a smaller dataset with every human’s predictions for all instances, or employing semi-supervised learning [Hemmer et al., 2023] to impute the missing human predictions. When deferring, human work capacity constraints are never considered in multi-expert L2D, where the goal is to find the best decision-maker for each instance, disregarding the amount of instances that are deferred to each individual agent. To the best of our knowledge, limitation of the amount of deferrals is only considered in the single-expert case, which allows for the inclusion of a regularization term to control the amount of deferrals [Mozannar and Sontag, 2020b, Steege, 2023].
@@
In our work, we propose a multi-expert L2D algorithm that can be used in cost-sensitive scenarios, trained with restrictive data requirements, and takes into account individual human work-capacity constraints that can be changed in testing.
@@ 2.2 Simulation of Human Experts
Due to the lack of sizeable, public, real-world datasets with multiple experts, most authors use label noise to produce arbitrarily accurate expert predictions on top of established datasets found in the ML literature. Mozannar and Sontag [2020b] use CIFAR-10 [Krizhevsky et al., 2009] and simulate an expert with perfect accuracy on a fraction of the 10 classes, but random accuracy on the others (see also the work by Verma and Nalisnick [2022b] and Charusaie et al. [2022]). The main drawback of these synthetic experts is that their expertise is rather simplistic, being either feature-independent or only dependent on a single feature or concept. This type of approach has been criticised [Zhu et al., 2021, Berthon et al., 2021], and instance-dependent label noise (IDN) has been proposed as a more realistic alternative, as human errors are likely to be dependent on the difficulty of a given task, and, as such, should depend on ithe features. In this work, we propose an IDN approach to simulate more complex and realistic synthetic experts.
@@ 3 Method
The main goal of our work is to develop a multi-expert assignment system that is able to optimize assignments in cost-sensitive tasks, subject to human work capacity constraints. For this method to be applicable in real-world scenarios, it is crucial that our assigner can be trained with limited human prediction data (only one expert prediction per instance) and that, in inference, our method be robust to variations in expert work capacity.
To tackle these objectives, we propose DeCCaF: a novel assignment system that optimizes instance allocation to a team of one or more analysts, while respecting their work capacity constraints. Given a set of instances with one associated expert prediction, we train a human expertise model (HEM) that jointly models the human team’s behavior. This model predicts the probability that deferral to a given expert will result in a correct decision. An ML classifier is trained over the same sample with the aim of estimating the likelihood of correctness in automatically predicting on a given instance. We then employ constraint programming (CP) to maximize the global probability of obtaining correct decisions under workload constraints.
@@ 3.1 Deferral Formulation
Assume that, for each instance $i$, there is a ground truth label $\mbox{y}_{i}\in\mathcal{Y}$, a vector $\bm{x}_{i}\in\mathcal{X}$ representing its features, a specific cost of misclassification $c_{i}$, and a prediction $m_{j,i}\in\mathcal{Y}$ from a given expert $j\in\{1,...,J\}$. The training set can then be represented as $S=\{\bm{x}_{i},\mbox{y}_{i},m_{j,i},c_{i}\}_{i=1}^{N}$.
@@
We first focus on building our L2D framework using the classifier-rejector approach [Madras et al., 2018a, Mozannar and Sontag, 2020a, Mozannar et al., 2023, Verma and Nalisnick, 2022a, Verma et al., 2023], first introduced by Cortes et al. [2016]. This approach focuses on learning two models: a classifier denoted as $h:\mathcal{X}\xrightarrow{}\mathcal{Y}$, and a rejector $r:\mathcal{X}\rightarrow\{0,1,...,J\}$. If $r(\bm{x}_{i})=0$, the classifier $h$ will make the decision on instance $i$; if $r(\bm{x}_{i})=j$, the decision on instance $i$ will be deferred to expert $j$. We will consider the $0-1$ loss proposed by Verma et al. [2023] as the learning objective. According to this formulation, when the classifier makes the prediction (i.e., $r(\bm{x})=0$), the system incurs a loss of 1 if the classifier is incorrect. When expert $j$ makes the prediction (i.e., $r(\bm{x})=j$), the system incurs a loss of 1 if the human is incorrect. Formally, the $0-1$ expected loss is defined as
$\begin{split}L_{0-1}=\mathbb{E}_{\bm{x},\mbox{\scriptsize y},\{m_{j}\}_{j=1}^{J}}&\Bigg{[}\mathbb{I}[r(\bm{x})=0]\;\mathbb{I}[h(\bm{x})\neq\mbox{y}]+\sum_{j=1}^{J}\mathbb{I}[r(\bm{x})=j]\;\mathbb{I}[m_{j}(\bm{x})\neq\mbox{y}]\Bigg{]}.\end{split}$ (1)
@@
Verma et al. [2023] demonstrate that the Bayes-optimal classifier and rejector, i.e., those that minimize $L_{0-1}$, satisfy
@@
$\begin{split}h^{B}(\bm{x})&=\operatorname*{arg\,max}_{\mbox{y}\in\mathcal{Y}}\mathbb{P}(\mbox{y}\lvert\bm{x})\\ r^{B}(\bm{x})&=\begin{cases}0\;\mbox{if}\;\mathbb{P}(\mbox{y}=h^{B}(\bm{x})\lvert\bm{x})>\mathbb{P}(\mbox{y}=m_{j}(\bm{x})\lvert\bm{x})\quad\forall j\\ \operatorname*{arg\,max}_{j\in\{1,...,J\}}\mathbb{P}(\mbox{y}=m_{j}(\bm{x})\lvert\bm{x})\quad\mbox{otherwise,}\end{cases}\end{split}$ (2)
@@
where $\mathbb{P}(\mbox{y}\lvert\bm{x})$ is the probability of the label under the data generating process, and $\mathbb{P}(\mbox{y}=m_{j}\lvert\bm{x})$ is the true probability that expert $j$ is be correct. As $L_{0-1}$ is non-convex, thus computationally hard to optimize, previous L2D work [Verma and Nalisnick, 2022a, Mozannar and Sontag, 2020a] focuses on the derivation of consistent convex surrogate losses, whose minimization results in $h^{*}$ and $r^{*}$ that approximate the Bayes optimal classifier-rejector pair. This approach assumes that $h$ and $r$ will be modelled by algorithms that fit the statistical query model [Kearns, 1998] (e.g. neural networks, decision trees), whose learning process involves the empirical approximation of the expected value of the surrogate losses by using a training sample $S$. The minimization of the approximation of the expected value (e.g. via gradient descent) produces a classifier $h^{*}$ and a rejector $r^{*}$ whose decisions are shown to converge to $h^{B}$ and $r^{B}$, given sufficient data.
@@
In this work, rather than deriving a consistent surrogate loss for simultaneously training $h$ and $r$, we train both components separately. In the following paragraphs, we propose a training process for the classifier and the rejector under limited data availability, demonstrating their convergence to $h^{B}$ and $r^{B}$. For simplicity, we consider the binary classification case, where $\mathcal{Y}=\{0,1\}$.

### v3

@@ Abstract
Learning to defer (L2D) aims to improve human-AI collaboration systems by learning how to defer decisions to humans when they are more likely to be correct than an ML classifier. Existing research in L2D overlooks key real-world aspects that impede its practical adoption, namely: i) neglecting cost-sensitive scenarios, where type I and type II errors have different costs; ii) requiring concurrent human predictions for every instance of the training dataset; and iii) not dealing with human work-capacity constraints. To address these issues, we propose the deferral under cost and capacity constraints framework (DeCCaF). DeCCaF is a novel L2D approach, employing supervised learning to model the probability of human error under less restrictive data requirements (only one expert prediction per instance) and using constraint programming to globally minimize the error cost, subject to workload limitations. We test DeCCaF in a series of cost-sensitive fraud detection scenarios with different teams of 9 synthetic fraud analysts, with individual work-capacity constraints. The results demonstrate that our approach performs significantly better than the baselines in a wide array of scenarios, achieving an average $8.4\%$ reduction in the misclassification cost. The code used for the experiments is available at https://github.com/feedzai/deccaf
1 Introduction
An increasing body of recent research has been dedicated to human-AI collaboration (HAIC), with several authors arguing that humans have complementary sets of strengths and weaknesses to those of AI (De-Arteaga et al., 2020; Dellermann et al., 2019). Collaborative systems have demonstrated that humans are able to rectify model predictions in specific instances (De-Arteaga et al., 2020), and have shown that humans collaborating with an ML model may achieve synergistic performance: a higher performance than humans or models alone (Inkpen et al., 2022). In high-stakes scenarios where ML models can outperform humans, such as healthcare (Gulshan et al., 2016), HAIC systems can help address safety concerns (e.g., the effect of changes in the data distribution (Gama et al., 2014)), by ensuring the involvement of humans in the decision-making process.
The state-of-the-art framework to manage assignments in HAIC is learning to defer (L2D) (Charusaie et al., 2022; Hemmer et al., 2022; Raghu et al., 2019b; a; Mozannar & Sontag, 2020b; Mozannar et al., 2023; Madras et al., 2018b; Steege, 2023; Verma & Nalisnick, 2022a; Verma et al., 2023). L2D aims to improve upon previous approaches, such as rejection learning (Chow, 1970; Cortes et al., 2016), which defer based solely on the ML model’s confidence, by also estimating the human confidence in a given prediction and passing the instance to the decision-maker who is most likely to make the correct decision.
Previous work in L2D does not address several key aspects of collaborative systems. In real-world scenarios, multiple human experts are employed to carry out the classification task, as the volume of instances to process cannot be handled by a single human. However, only a small subset of L2D research focuses on the multi-expert setting, where decisions have to be distributed throughout a team comprised of a single ML classifier and multiple human experts (Keswani et al., 2021; Hemmer et al., 2022; Verma et al., 2023). To the best of our knowledge, Verma et al. (2023) propose the only two consistent multi-expert L2D formulations, with both approaches assuming the existence of every expert’s predictions for all training instances. In real-world applications, we often have more limited data availability, with only a single expert (De-Arteaga et al., 2020) or a small subset of the experts (Gulshan et al., 2016) providing predictions for each instance. This means that a practitioner aiming to train L2D algorithms would have to purposefully collect the set of every expert’s predictions for a sample of instances, which could incur large costs or even be unfeasible. We will propose L2D architectures that allow for assigners to be trained assuming that each instance is accompanied by the prediction of only one expert out of the team.
Current multi-expert L2D methods also neglect human capacity limitations. Should there be an expert that consistently outperforms the model and all the other experts, the optimal assignment would be to defer all cases to that expert, which in practice is not feasible. Furthermore, current work neglects cost-sensitive scenarios, where the cost of misclassification can be class or even instance-dependent (e.g., in medicine, false alarms are generally considered less harmful than failing to diagnose a disease).
@@ 1 Introduction
To address the aforementioned L2D limitations, we propose the deferral under cost and capacity constraints framework (DeCCaF): a novel deferral approach to manage assignments in cost-sensitive human-AI decision-making, which respects human capacity constraints. Our method is comprised of three separate components, represented schematically in Figure 1: (1) an ML classifier modelling the probability of the target class given the instance features; (2) a human expertise model (HEM) that models the probabilities of correctness of each of the experts in the team; and (3) an assigner that computes the best possible set of assignments given misclassification costs and capacity constraints.
Due to the lack of sizeable datasets with multiple human predictions and to the high costs associated with producing one, we empirically validate DeCCaF in a series of realistic cost-sensitive fraud detection scenarios, where a team of 9 synthetic fraud analysts and one ML classifier are tasked with reviewing financial fraud alerts. We conclude that, across all scenarios, DeCCaF performs similarly to, or significantly better than L2D baselines, as measured by average misclassification costs.
@@
In Section 2, we describe the relevant related work, focusing on recent developments in multi-expert L2D and examining the shortcomings of the deferral systems proposed so far. We also discuss current practices in L2D evaluation, particularly the use of synthetically generated expert predictions, and how they can be improved. We then describe DeCCaF in Section 3, first by formulating a novel training method, compatible with limited data availability, for the commonly used classifier-rejector framework (Mozannar & Sontag, 2020a; Verma & Nalisnick, 2022a; Verma et al., 2023; Cortes et al., 2016), demonstrating that it produces optimal unconstrained deferral. As the traditional classifier-rejector approach fails to consider the existence of human work-capacity limitations, we propose a novel formulation for global assignment optimization under capacity constraints. In Section 4 we detail a realistic fraud detection experimental setup as well as the method for the synthetic expert prediction generation. We also discuss the capacity-aware baselines used in our L2D benchmark. The experimental results are reported and analyzed in Section 5.
@@ 2.1 Current L2D Methods
The simplest deferral approach in the literature is rejection learning (ReL), which dates back to the work of Chow (1970); Cortes et al. (2016). In a HAIC setting, ReL defers to humans the instances that the model rejects to predict (Madras et al., 2018b; Raghu et al., 2019a). A simple example (Hendrycks & Gimpel, 2017) is to obtain uncertainty estimates of the model prediction for each instance, rejecting to predict if the uncertainty estimate is above a given threshold.
@@
Madras et al. (2018b) criticize ReL, arguing that it does not consider the performance of the human involved in the task and propose learning to defer (L2D), where the classifier and assignment system are jointly trained, taking into account a single model and a single human and accounting for human error in the training loss. Many authors have since contributed to the single-expert framework: Mozannar & Sontag (2020b) show that the loss proposed by Madras et al. (2018b) is inconsistent, proposing a consistent surrogate loss that yields better results in testing; Verma & Nalisnick (2022b) critique the approach of Mozannar & Sontag (2020b), demonstrating that their surrogate loss has a degenerate parameterization, causing miscalibration of the estimated probability of expert correctness. Keswani et al. (2021) observe that decisions can often be deferred to one or more humans out of a team, expanding L2D to the multi-expert setting. Verma et al. (2023) propose the first consistent and calibrated surrogate losses for the multi-expert setting, by adapting the work of Verma & Nalisnick (2022b) and the softmax surrogate loss of Mozannar & Sontag (2020b). All aforementioned studies focus on deriving surrogates for the $0-1$ loss, meaning they are not directly applicable to cost-sensitive scenarios, where the cost of erring can be class-dependent (i.e., different costs for false positive and false negative errors), or even instance-specific (i.e., every instance has an associated misclassification cost).
@@
Another key facet of L2D research is the interplay between the two components of an assignment system: the rejector (which decides whether to defer and to whom) and the classifier (which produces automatic predictions if the rejector chooses not to defer). Mozannar & Sontag (2020b) argue that the main classifier should specialize on the instances that will be assigned to it, in detriment of those that will not. This is done by jointly training the classifier and the rejector, without penalizing the classifier’s mistakes on instances that the rejector defers to an expert. The approach of Verma et al. (2023) differs in that the rejector and the classifier are trained independently, meaning that the classifier is encouraged to predict correctly on all instances, likely to be deferred or not. However, in the multi-expert setting, the two-stage one-vs.-all approach of Verma et al. (2023) outperforms their adaptation of the softmax loss of Mozannar & Sontag (2020a), which employs joint-learning, due to the same calibration problems observed in the single-expert setting. A more recent joint learning approach proposed by Mozannar et al. (2023) is theoretically shown to enable the system to achieve an improvement in overall performance, as their surrogate loss differs from previous work in that it is realizable $\mathcal{H}$-consistent. In more recent work, however, Mao et al. (2024) demonstrate that two-staged learning, the separate training of the classifier and rejector, with their proposed surrogate losses, can also guarantee realizable consistency. We argue that joint learning is not be suitable for real-world applications, as, by design, the instances that are most likely to be deferred are those in which the classifier will perform worse, which would make the system highly susceptible to changes in human availability, should the AI have to predict on instances that were originally meant to be deferred.
@@
As previously mentioned, the key barrier to the adoption of current L2D methods is that they require predictions from every human in the team, for every training instance. Current L2D research tackling data-efficient learning focuses only on the single-expert setting, using active-learning principles (Charusaie et al., 2022) to construct a smaller dataset with every human’s predictions for all instances, or employing semi-supervised learning (Hemmer et al., 2023) to impute the missing human predictions.
When deferring, constraints in human work-capacity are rarely considered in multi-expert L2D, where the goal is usually to find the best decision-maker for each instance, disregarding the amount of instances that are deferred to each individual agent. To the best of our knowledge, only Mao et al. (2024) consider control of the amount of deferrals in multi-expert L2D, by allowing for the definition of a constant deferral cost $c_{j}$ for each available expert $j$, which is included in their realizable $\mathcal{H}$-consistent two stage losses. Similar approaches had been considered in the single-expert case, allowing for the inclusion of a regularization term or expert deferral cost (Mozannar & Sontag, 2020b; Steege, 2023; Narasimhan et al., 2022). While such a formulation allows for regularization of deferral across experts, it would not be adequate for a real-world system, where the availability of experts may vary drastically over time. In a real-world system, different teams of experts may work at different schedules or have days off, which would result in experts being available during certain time intervals while absent in others.
The existence of changes in expert availability within real-world HAIC scenarios means that an L2D system must be able to accommodate expert capacity constraints which are defined at inference-time, and not during training. This restriction is similar to that considered by Zhou et al. (2022) in the context of sparsely activated mixture-of-experts models, where the routing strategy (i.e., how tokens are distributed throughout experts) is akin to the task of L2D algorithms. The authors propose allowing for a hard constraint on the number of tokens selected by each expert and the number of experts each token is routed to in inference. This approach allows for dynamic limitation of the compute resources necessary, without needing to retrain their models.
@@
In our work, we propose a multi-expert L2D algorithm that can be used in cost-sensitive scenarios, trained with restrictive data requirements, and taking into account individual human work-capacity constraints that are imposed only at inference time.
@@ 2.2 Simulation of Human Experts
Due to the lack of sizeable, public, real-world datasets with multiple experts, most authors use label noise to produce arbitrarily accurate expert predictions on top of established datasets found in the ML literature. Mozannar & Sontag (2020b) use CIFAR-10 (Krizhevsky et al., 2009) and simulate an expert with perfect accuracy on a fraction of the 10 classes, but random accuracy on the others (see also the work by Verma & Nalisnick (2022b) and Charusaie et al. (2022)). The main drawback of these synthetic experts is that their expertise is rather simplistic, being either feature-independent or only dependent on a single feature or concept. This type of approach has been criticised (Zhu et al., 2021; Berthon et al., 2021), and instance-dependent label noise (IDN) has been proposed as a more realistic alternative, as human errors are likely to be dependent on the difficulty of a given task, and, as such, should depend on its features. In this work, we propose an IDN approach to simulate more complex and realistic synthetic experts.
@@ 3 Method
The main goal of our work is to develop a multi-expert assignment system that is able to optimize assignments in cost-sensitive tasks, subject to human work-capacity constraints. For this method to be applicable in real-world scenarios, it is crucial that the assigner can be trained with limited human prediction data (only one expert prediction per instance) and that, in inference, our method be robust to variations in expert work-capacity.
To tackle these objectives, we propose DeCCaF: a novel assignment system that optimizes instance allocation to a team of one or more analysts, while respecting their work-capacity constraints. Given a set of instances with one associated expert prediction, we train a human expertise model (HEM) that jointly models the human team’s behavior. This model predicts the probability that deferral to a given expert will result in a correct decision. An ML classifier is trained over the same sample with the aim of estimating the likelihood of correctness in automatically predicting on a given instance. We then employ constraint programming (CP) to maximize the global probability of obtaining correct decisions under workload constraints.
@@ 3.1 Deferral Formulation
Assume that, for each instance $i$, there is a ground truth label $y_{i}\in\mathcal{Y}$, a vector $\bm{x}_{i}\in\mathcal{X}$ representing its features, a specific cost of misclassification $c_{i}$, a prediction $m_{j,i}\in\mathcal{Y}$ from a given expert $j\in\{1,...,J\}$, and $j_{i}$, identifying which expert made the prediction. The training set can then be represented as $S=\{\bm{x}_{i},y_{i},m_{j,i},c_{i},j_{i}\}_{i=1}^{N}$.
@@
We first focus on building our L2D framework using the classifier-rejector approach (Madras et al., 2018a; Mozannar & Sontag, 2020a; Mozannar et al., 2023; Verma & Nalisnick, 2022a; Verma et al., 2023), first introduced by Cortes et al. (2016). This approach focuses on learning two models: a classifier denoted as $h:\mathcal{X}\xrightarrow{}\mathcal{Y}$ and a rejector $r:\mathcal{X}\rightarrow\{0,1,...,J\}$. If $r(\bm{x}_{i})=0$, the classifier $h$ will make the decision on instance $i$; if $r(\bm{x}_{i})=j$, the decision on instance $i$ will be deferred to expert $j$. We will consider the $0-1$ loss, as proposed by Verma et al. (2023), as the learning objective. According to this formulation, when the classifier makes the prediction (i.e., $r(\bm{x})=0$), the system incurs a loss of 1 if the classifier is incorrect. When expert $j$ makes the prediction (i.e., $r(\bm{x})=j$), the system incurs a loss of 1 if the human is incorrect. Formally, the $0-1$ expected loss is defined as
$\begin{split}L_{0-1}(h,r)=\mathbb{E}_{\bm{x},\mbox{\scriptsize y},\{m_{j}\}_{j=1}^{J}}&\Bigg{[}\mathbb{I}[r(\bm{x})=0]\;\mathbb{I}[h(\bm{x})\neq y]+\sum_{j=1}^{J}\mathbb{I}[r(\bm{x})=j]\;\mathbb{I}[m_{j}(\bm{x})\neq y]\Bigg{]}.\end{split}$ (1)
@@
Verma et al. (2023) demonstrate that the Bayes-optimal classifier and rejector, i.e., those that minimize $L_{0-1}$, satisfy
$h^{*}(\bm{x})=\displaystyle\operatorname*{arg\,max}_{y\in\mathcal{Y}}\mathbb{P}(y\lvert\bm{x})$ (2)
@@
$r^{*}(\bm{x})=\begin{cases}0\;&\mbox{if}\;\mathbb{P}(y=h^{*}(\bm{x})\lvert\bm{x})>\mathbb{P}(y=m_{j}(\bm{x})\lvert\bm{x}),\;\forall j\in\{1,...,J\}\\ \displaystyle\operatorname*{arg\,max}_{j\in\{1,...,J\}}\mathbb{P}(y=m_{j}(\bm{x})\lvert\bm{x})\;&\mbox{otherwise,}\end{cases}$ (3)
@@
where $\mathbb{P}(y\lvert\bm{x})$ is the probability of the label under the data generating process, and $\mathbb{P}(y=m_{j}\lvert\bm{x})$ is the true probability that expert $j$ is be correct. As $L_{0-1}$ is non-convex, thus computationally hard to optimize, previous L2D work (Verma & Nalisnick, 2022a; Mozannar & Sontag, 2020a) focuses on the derivation of consistent convex surrogate losses, whose minimization results in $\hat{h}$ and $\hat{r}$ that approximate the Bayes optimal classifier-rejector pair. This approach assumes that $h$ and $r$ will be modelled by algorithms that fit the statistical query model (Kearns, 1998) (e.g., neural networks, decision trees), whose learning process involves the empirical approximation of the expected value of the surrogate losses by using a training sample $S$. The minimization of the approximation of the expected value (e.g., via gradient descent) produces a classifier $\hat{h}$ and a rejector $\hat{r}$ whose decisions are shown to converge to $h^{*}$ and $r^{*}$ as the amount of data increases.
@@
In this work, rather than deriving a consistent surrogate loss for simultaneously training $h$ and $r$, we train both components separately. In the following paragraphs, we propose a training process for the classifier and the rejector under limited data availability, demonstrating their convergence to $h^{*}$ and $r^{*}$. For simplicity, we consider the binary classification case, where $\mathcal{Y}=\{0,1\}$.